Solve using either substitution or addition method. Tell which method was used?
3y=x+5
-3x+8y=8

Answer by checkley77(12844)

3y=x+5 multiply by -3 & add.
-3x+8y=8
3x-9y=-15
-----------
-y=-7
y=7 ans.
3*7=x+5
21=x+5
x=-21-5
x=-26 ans.
Proof:
-3*-26-9*7=-15
78-63=-15
-15=-15